OpenStudy (anonymous):

If g(x) = 3x + 1, evaluate g(-14). a.-41 b.-10 c. -5

OpenStudy (anonymous):

a

OpenStudy (anonymous):

3(-14) +1 = -42 +1 = -41

OpenStudy (callisto):

g(-14) = 3(-14) + 1= -42+1 = -41